# Colloquium: Synthetic quantum matter in non-standard geometries

### v1

@@ I Introduction and outline
Quasicrystals | Fractals | 4D Space | Curved Space | |
Atoms | Optical Superlattice 1D, 2D Quasicrystal Optical Lattices, Momentum-Space Lattices | Optical Tweezers | Synthetic Dimension, Charge Pumping in 2D Optical Superlattice | Speed of Sound Variation in Inhomogeneous BECs Expanding Bubble Traps |
Photons/QED | Modulated Micro-wires, 1D Lattice of Micro-pillars, Waveguide Arrays, Superconducting Qubits | Waveguide Arrays | Charge Pumping in 2D Wave Guide Arrays | Superconducting Qubits in Hyperbolic Lattice, Strained Photonic Graphene Lattice |
Electrons | Synthetic Penrose Tiling with Electron on Surface | Molecular Assembly, Scanning Tunneling Microscope | 4D Lattice Embedded in 3D Electric Circuit | Hyperbolic Lattice Electric Circuit |
@@
Quasicrystals are a third kind of solids, between the periodic crystals and the amorphous solids or glasses, which, typically due to a non-negligible amount of disorder, lack translational symmetry. The structure of quasicrystals follows a deterministic construction rule, which gives rise to long-range order as in ordinary crystals. However, like glasses, they are aperiodic. Physically, their order becomes manifest, for instance, through Bragg peaks in a diffraction experiment, which also has played a crucial role in the discovery of quasicrystals in 1984 [221, 141], when D. Shechtman and co-workers found aluminum-manganese alloys with tenfold rotational symmetry [221]. Indeed, some definitions of a quasicrystal [147] demand that it exhibits a rotational symmetry incompatible with periodic structures limited to two-, three-, four-, or sixfold rotational symmetry. Shechtman’s discovery awarded the Nobel Prize in Chemistry in 2011, extended the research on aperiodic tilings from mathematically and/or aesthetically grounded studies in the 1970s [186] to the fields of physics and material science. Nowadays, research on quasicrystals is gaining importance through the rapid progress of synthetic materials, which have opened the door to experimental studies of an even larger variety of quasicrystals, including paradigmatic 1D models. On the tight-binding level, these models are of the form
with () creation (annihilation) operators on site . The parameters and are neither periodic (as in crystal models) nor random (as in glassy models), but (at least) one of them is obtained from a quasiperiodic sequence, such as the Fibonacci sequence, see Jagannathan [116] for a review. For instance, the diagonal Fibonacci quasicrystal has constant hopping, , but the potential is a sequence of binary potential values. This sequence is generated iteratively by merging the previous two iterations of the sequence. Another famous 1D quasicrystal model is the Aubry-André (AA) model with
@@
An example of a fractal, the so-called Sierpiński gasket, is shown on the top of the second column of Table 1. The structure is generated by dividing an equilateral triangle into smaller ones and by removing the central smaller triangle. The Hausdorff dimension of this structure is . A similar structure, starting from a square instead of a triangle, is known as Sierpiński carpet and has a Hausdorff dimension of . The procedure can also be applied to structures embedded in a three-dimensional space, e.g., tetrahedrons. By division into smaller and smaller tetrahedrons combined with the removal of selected tetrahedrons, one ends up at the Sierpiński tetrahedron, which curiously has an integer Hausdorff dimension of 2. Through synthetic systems, fractals have also entered the arena of quantum physics, see Table 1, and in particular, the dynamics of quantum particles in fractal space have become the subject of intense research.
@@
Finally, lattice systems can become exotic if their metric is not the usual Euclidean one but exhibits some curvature. Curved spaces have been traditionally investigated in high-energy physics and cosmology. There has been a growing interest in realizing tabletop simulations of hyperbolic lattices in recent years. The motivations arise from discovering the holographic principle [158, 277] and the characterization of classical and quantum states in spaces with negative curvature. One possible way to investigate curved lattices beyond the Euclidean case is to work on the Poincaré disk. Here, the tessellation of the space is defined via two indices , known as the Schläfli symbol. It corresponds to a tessellation of the plane by regular -gons with coordination number [59]. Given a , the curvature of the space is defined by
If , the space is spherical (positive curvature); whereas, if , the space is Euclidean (zero curvature), and finally, if , we have a hyperbolic space (negative curvature) [37]. Euclidean tesselations are limited to the well-known triangular , square , and hexagonal lattices. The example of a hyperbolic lattice is shown in the illustration of Table 1 (right column). Hence, by the choice of connectivity in a lattice , non-Euclidean spaces can be implemented, and in this way hyperbolic lattices, characterized by negative curvature, have been produced in photonic and electronic quantum simulators, see Table 1. Another approach to curved spaces is through an analogy between sound propagation on a background hydrodynamic flow and field propagation in curved space-time, see Unruh [250]. This analogy has allowed for experimental studies of gravitational phenomena in flat quantum simulators.
This colloquium reviews how the described variety of exotic geometries is implemented and exploited in contemporary quantum simulators using atomic, photonic, and electronic platforms. Specifically, in the next section (Implementation, Sec. II), we will review state-of-the-art quantum engineering techniques used to achieve such geometries. In the subsequent section (Applications, Sec. III), we will discuss the specific implications of exotic geometries on physical system behavior. Here, we will differentiate between three different phenomenological areas: (i) localization phenomena, (ii) topological phenomena, (iii) analog black holes and cosmology.
@@ II.1 Atomic systems
Ultracold atoms are a formidable experimental platform for quantum simulation due to their large tunability of confining potentials and interaction strength, and their ability to reach strongly-correlated quantum states with long lifetimes. Neutral atoms are routinely cooled to Mikrokelvin temperatures via laser cooling and further cooled to Nanokelvin temperatures via evaporative cooling in magnetic or optical traps. Ultrahigh vacuum conditions allow for thermal insulation from the vacuum chamber walls. Atoms exist as bosons or fermions depending on the total spin of the atomic isotope, and at Nanokelvin temperatures, they form Bose-Einstein condensates (BEC) or degenerate Fermi gases, respectively.
@@
Atoms have internal hyperfine states that can equally be controlled and can be harnessed as an effective spin. This allows to study quantum magnetism or to engineer spin-orbit coupling via two-photon Raman transitions. Some atoms also have long-lived optically excited states that can be employed as orbital degrees of freedom accessed by single-photon transitions.
@@ II.1.1 Quasicrystals and fractals
Optical lattices have been a primary tool for quantum simulation with cold atoms, including exotic geometries [276]. Very elegantly, optical lattices can also produce quasicrystalline potentials by an appropriate choice of laser beams forming the lattice. The AA model in Eq. (2), a 1D lattice with quasi-random on-site energies, was realized by adding a second weaker 1D optical lattice of incommensurate spatial period due to an incommensurate wavelength [200, 197]. This approach can be used as an alternative to laser speckle disorder [31], but as will be discussed in Sec. III.1, also allows for fine-tuned localization properties, such as 1D mobility edges. In 2D, fivefold quasicrystal potentials would be formed by the interference of five lattice beams [211], and a challenge would be to control the relative phases of the laser beams [58]. This technical challenge can be overcome using the multi-frequency lattice scheme leading to a pairwise interference of the beams at different frequencies with well-controlled relative phases [129]. An eightfold quasicrystal potential was realized by superposing four 1D optical lattices at slightly different frequencies such that they do not interfere with each other [255, 212] — see Figs. 1a,b. These experiments have also triggered new theoretical interest in many-body physics in quasicrystal potentials and new descriptions such as Hubbard models for quasicrystalline potentials [118, 94, 292].
@@ II.1.3 Curved space
An exciting way of studying an expanding universe is to work with BECs in supersonically expanding toroidal traps [28]. This was realized with 1D ring traps with variable radius created via a digital micromirror device [72, 17], finding a redshift of long-wavelength excitations as in an expanding universe as well as the production of solitons and vortices. A 2D version, a bubble trap forming the surface of a sphere [91], was recently realized via radiofrequency dressing of a magnetic trap using the microgravity conditions on the international space station [45]. The experiment studied the thermodynamics of ultracold bosonic atoms, demonstrating substantial cooling upon increasing the size of the sphere, which links to expansion physics analogue to the expanding universe. Such 2D bubble traps are also relevant for the study of space-curvature effects on quantum many-body physics, e.g., on superfluidity and the dynamics of vortices and solitons [42, 248, 247].
@@ II.2.2 Lattices with synthetic dimensions
Interestingly, frequency and time-bin, both related to the temporal aspect of space-time, are two of the most commonly used degrees of freedom for synthetic dimensions in topological photonics. Discrete resonances in ring resonators, shown in Fig. 5a, provide a frequency lattice where nearest-neighbor couplings can be engineered on-demand using electro-optical modulation [182, 287]. This configuration has been implemented in various material platforms, successfully emulating different facets of topological physics [70, 69, 264, 16, 185]. Regarding using time-bin as a synthetic dimension, the scheme involves studying light propagation in two loops of optical fiber of different lengths connected by a beam-splitter. This setup makes the system equivalent to a 2D lattice, where one dimension corresponds to the location of the quantum walk and the other to the time step [275, 274, 269], as shown in Figure 5(b). Additionally, tuning the beamsplitter behavior over time can introduce another synthetic dimension to explore new topological phases, such as the anomalous Floquet metal [2].
@@
Synthetic momenta, which differ from synthetic dimensions in spatial space, have recently emerged as a promising concept for exploring higher-dimensional topology. The key idea is to use spatial quantities as synthetic momenta, thereby naturally breaking time-reversal symmetry in the extended momentum space. For instance, in misaligned photonic gratings that share the same period and are evanescently coupled, the relative shifts between the gratings serve as synthetic momenta [175, 137, 174], as shown in Fig. 5c. Using this concept, chiral edge state [175] and 2D nodal lines [137] were proposed for (1+1)-dimensional systems of bilayer gratings, while Weyl semimetal physics with Fermi arc reconstruction [174] was proposed for (1+2)-dimensional systems of trilayer gratings. Given the vibrant field of multilayer photonics and various recent realizations [241, 240], the experimental implementation of these synthetic momenta systems in the near future is anticipated, potentially addressing 4D or 5D physics.
@@ II.3 Electronic systems
The last platform we will consider in this section consists of creating artificial lattices in electronic-based systems. We will mainly introduce four different approaches in electronic platforms: (i) electrons on surface, (ii) molecular self-assembly, (iii) classical electric circuits, and (iv) circuit quantum electrodynamics (QED). In these cases, the goal is to simulate a specific lattice model. At present, the first three platforms are limited to simulated single-particle physics. In the first case, the electrons of a 2DEG are forced to a specific potential with a specific geometry. Usually, the Shockley surface states of metals are considered for the 2DEG [223], e.g., Cu, or Ag on the (111) surface. The lattice potential is created by placing atoms or molecules with atomic precision using the tip of a scanning tunneling microscope (STM). The tip of the STM is successively used for measuring the single-particle local density of states (LDOS) at the different positions by the differential conductance of the system as a function of the applied tip-substrate bias . This quantity is proportional to the modulus square of the system wave function at a given position and voltage. The major drawback of this technique is the finite lifetime of the surface electronic states. Adatoms of the surface scatter surface electrons into the bulk of the system, leading to the broadening of the electronic resonances. The broadening at 4.6 K is generally smaller in Ag compared to Cu [75]. However, the adatoms placed on Cu, especially CO molecules, are more stable and easier to move on Cu than Ag, making Cu the metallic reference system for creating artificial lattice structures [85]. This combination has been employed for designing lattices with translational invariance [93, 226, 225, 119, 84, 145].
In the case of the molecular self-assembly approach, molecular precursors are deposited on a metallic surface; after annealing, the precursors rearrange themself in structures that can be more complex than the original precursors, either by following a specific pattern intrinsic to the precursor or because of some property of the substrate. Successively, the structure is analyzed either by STM or by an atomic force microscope [189].
@@ II.3.3 Higher dimensional physics
Classical electric circuits are also a versatile platform for implementing higher-dimensional physics. The two-dimensional array of capacitors and inductors can be easily extended to realize periodic 3D and 4D circuits on a breadboard [291]. The idea is to decompose an D structure into a set of D circuit boards connected to each other, see Fig. 8.

### v2

@@ I Introduction and outline
Quasicrystals are a third kind of solids, between the periodic crystals and the amorphous solids or glasses, which, typically due to a non-negligible amount of disorder, lack translational symmetry. The structure of quasicrystals follows a deterministic construction rule, which gives rise to long-range order as in ordinary crystals. However, like glasses, they are aperiodic. Physically, their order becomes manifest, for instance, through Bragg peaks in a diffraction experiment, which also has played a crucial role in the discovery of quasicrystals in 1984 [237, 154], when D. Shechtman and co-workers found aluminum-manganese alloys with tenfold rotational symmetry [237]. Indeed, some definitions of a quasicrystal demand that it exhibits a rotational symmetry incompatible with periodic structures limited to two-, three-, four-, or sixfold rotational symmetry. According to a more general definition that also admits 1D quasicrystals, this term is simply an abbreviation for quasiperiodic crystal [162].
Shechtman’s discovery, awarded the Nobel Prize in Chemistry in 2011, extended the research on aperiodic tilings from mathematically and/or aesthetically grounded studies in the 1970s [202] to physics and material science. Nowadays, research on quasicrystals is gaining importance through the rapid progress of synthetic materials, which have opened the door to experimental studies of an even larger variety of quasicrystals, including paradigmatic 1D models. On the tight-binding level, these models are of the form
with () creation (annihilation) operators on site . The parameters and are neither periodic (as in crystal models) nor random (as in glassy models), but (at least) one of them is obtained from a quasiperiodic sequence, such as the Fibonacci sequence, see Jagannathan [126] for a review. For instance, the diagonal Fibonacci quasicrystal has constant hopping, , but the potential is a sequence of binary potential values. This sequence is generated iteratively by merging the previous two iterations. Another famous 1D quasicrystal model is the Aubry-André (AA) model with
@@
An example of a fractal, the so-called Sierpiński triangle or gasket, is shown on the top of the second column of Table 1. The structure is generated by dividing an equilateral triangle into smaller ones and removing the central smaller triangle. The Hausdorff dimension of this structure is . A similar structure, starting from a square instead of a triangle, is known as Sierpiński carpet and has a Hausdorff dimension of . The procedure can also be applied to structures embedded in a three-dimensional space, e.g., tetrahedrons. By division into smaller and smaller tetrahedrons combined with the removal of selected tetrahedrons, one ends up at the Sierpiński tetrahedron, which curiously has an integer Hausdorff dimension of 2. Through synthetic systems, fractals have also entered the arena of quantum physics, see Table 1, and in particular, the dynamics of quantum particles in fractal space have become the subject of intense research.
@@
Finally, lattice systems can become exotic if their metric is not the usual Euclidean but exhibits some curvature. Curved spaces have been traditionally investigated in high-energy physics and cosmology. In recent years, there has been a growing interest in realizing tabletop simulations of hyperbolic lattices. The motivations arise from discovering the holographic principle [174, 299] and the characterization of classical and quantum states in spaces with negative curvature. One possible way to investigate curved lattices beyond the Euclidean case is to work on the Poincaré disk. Here, the tessellation of the space is defined via two indices , known as the Schläfli symbol. It corresponds to a tessellation of the plane by regular -gons with coordination number [62]. Given a , the curvature of the space is defined by
If , the space is spherical (positive curvature); whereas, if , the space is Euclidean (zero curvature), and finally, if , we have a hyperbolic space (negative curvature) [38]. Euclidean tesselations are limited to the well-known triangular , square , and hexagonal lattices. The example of a hyperbolic lattice is shown in the illustration of Table 1 (right column). Hence, by the choice of connectivity in a lattice , non-Euclidean spaces can be implemented, and in this way, hyperbolic lattices, characterized by negative curvature, have been produced in photonic and electronic quantum simulators, see Table 1. Another approach to curved spaces is through an analogy between sound propagation on a background hydrodynamic flow and field propagation in curved space-time, see Unruh [270]. This analogy has allowed for experimental studies of gravitational phenomena in flat quantum simulators.
This colloquium reviews how the described variety of exotic geometries is implemented and exploited in contemporary quantum simulators using atomic, photonic, and electronic platforms. Specifically, in the next section (Implementation, Sec. II), we will review state-of-the-art quantum engineering techniques used to achieve such geometries. In the subsequent section (Applications, Sec. III), we will discuss the specific implications of exotic geometries on physical system behavior. Here, we will differentiate between three different phenomenological areas: (i) localization phenomena, (ii) topological phenomena, and (iii) analog black holes and cosmology.
@@ II.1 Atomic systems
Ultracold atoms are a formidable experimental platform for quantum simulation due to their large tunability of confining potentials and interaction strength and ability to reach strongly-correlated quantum states with long lifetimes. Neutral atoms are routinely cooled to Mikrokelvin temperatures via laser cooling and further cooled to Nanokelvin temperatures via evaporative cooling in magnetic or optical traps. Ultrahigh vacuum conditions allow for thermal insulation from the vacuum chamber walls. Atoms exist as bosons or fermions depending on the total spin of the atomic isotope, and at Nanokelvin temperatures, they form Bose-Einstein condensates (BEC) or degenerate Fermi gases, respectively.
@@
Atoms have internal hyperfine states that can equally be controlled and harnessed as an effective spin. This allows to study quantum magnetism or to engineer spin-orbit coupling via two-photon Raman transitions. Some atoms also have long-lived optically excited states that can be employed as orbital degrees of freedom accessed by single-photon transitions.
@@ II.1.1 Quasicrystals and fractals
Optical lattices have been a primary tool for quantum simulation with cold atoms, including exotic geometries [298]. Very elegantly, optical lattices can also produce quasicrystalline potentials by an appropriate choice of laser beams forming the lattice. The AA model in Eq. (2), a 1D lattice with quasi-random on-site energies, was realized by adding a second weaker 1D optical lattice of incommensurate spatial period due to an incommensurate wavelength [216, 213]. This approach can be used as an alternative to laser speckle disorder [32], but as will be discussed in Sec. III.1 also allows for fine-tuned localization properties, such as 1D mobility edges. In 2D, fivefold quasicrystal potentials would be formed by the interference of five lattice beams [227], and a challenge would be to control the relative phases of the laser beams [61]. This technical challenge can be overcome using the multi-frequency lattice scheme leading to a pairwise interference of the beams at different frequencies with well-controlled relative phases [141]. An eightfold quasicrystal potential was realized by superposing four 1D optical lattices at slightly different frequencies such that they do not interfere with each other [276, 228] — see Figs. 1a,b. These experiments have also triggered new theoretical interest in many-body physics in quasicrystal potentials and new descriptions such as Hubbard models for quasicrystalline potentials [128, 100, 103, 317].
@@ II.1.3 Curved space
An exciting way of studying an expanding universe is to work with BECs in supersonically expanding toroidal traps [29]. This was realized with 1D ring traps with variable radius created via a digital micromirror device [78, 18], finding a redshift of long-wavelength excitations as in an expanding universe as well as the production of solitons and vortices. A 2D version, a bubble trap forming the surface of a sphere [99], was recently realized via radiofrequency dressing of a magnetic trap using the microgravity conditions on the international space station [47]. The experiment studied the thermodynamics of ultracold bosonic atoms, demonstrating substantial cooling upon increasing the size of the sphere, which links to expansion physics analogue to the expanding universe. Such 2D bubble traps are also relevant for the study of space-curvature effects on quantum many-body physics, e.g., on superfluidity and the dynamics of vortices and solitons [44, 267, 168, 268, 266].
@@ II.2.2 Lattices with synthetic dimensions
Interestingly, frequency and time-bin, both related to the temporal aspect of space-time, are two of the most commonly used degrees of freedom for synthetic dimensions in topological photonics. Discrete resonances in ring resonators, shown in Fig. 5a, provide a frequency lattice where nearest-neighbor couplings can be engineered on-demand using electro-optical modulation [198, 310]. This configuration has been implemented in various material platforms, successfully emulating different facets of topological physics [75, 74, 285, 17, 201]. Regarding using time-bin as a synthetic dimension, the scheme involves studying light propagation in two loops of optical fiber of different lengths connected by a beam-splitter. This setup makes the system equivalent to a 2D lattice, where one dimension corresponds to the location of the quantum walk and the other to the time step [297, 296, 291], as shown in Fig. 5(b). Additionally, tuning the beamsplitter behavior over time can introduce another synthetic dimension to explore new topological phases, such as the anomalous Floquet metal [2].
@@
Synthetic momenta, which differ from synthetic dimensions in spatial space, have recently emerged as a promising concept for exploring higher-dimensional topology. The key idea is to use spatial quantities as synthetic momenta, thereby naturally breaking time-reversal symmetry in the extended momentum space. For instance, in misaligned photonic gratings that share the same period and are evanescently coupled, the relative shifts between the gratings serve as synthetic momenta [191, 150, 190], as shown in Fig. 5c. Using this concept, chiral edge state [191] and 2D nodal lines [150] were proposed for (1+1)-dimensional systems of bilayer gratings, while Weyl semimetal physics with Fermi arc reconstruction [190] was proposed for (1+2)-dimensional systems of trilayer gratings. Recently, synthetic momentum has been employed to explore non-orientable manifolds to circumvent the Nielsen–Ninomiya theorem [87]. Given the vibrant field of multilayer photonics and various recent realizations [259, 258], the experimental implementation of these synthetic momenta systems in the near future is anticipated, potentially addressing 4D or 5D physics.
@@ II.3 Electronic systems
The last platform we will consider in this section consists of creating artificial lattices in electronic-based systems. We will mainly introduce five different approaches in electronic platforms: (i) electrons on surface, (ii) molecular self-assembly, (iii) twisted layers of two-dimensional materials, (iv) classical electric circuits, and (v) circuit quantum electrodynamics (QED). In these cases, the goal is to simulate a specific lattice model. Some of these platforms are limited to simulate single-particle physics.
In case (i) of electrons on a surface, the electrons of a 2DEG are forced to a specific potential with a specific geometry. Usually, the Shockley surface states of metals are considered for the 2DEG [239], e.g., Cu, or Ag on the (111) surface. The lattice potential is created by placing atoms or molecules with atomic precision using the tip of a scanning tunneling microscope (STM). The tip of the STM is successively used for measuring the single-particle local density of states (LDOS) at the different positions by the differential conductance of the system as a function of the applied tip-substrate bias . This quantity is proportional to the modulus square of the system wave function at a given position and voltage. The major drawback of this technique is the finite lifetime of the surface electronic states. Adatoms of the surface scatter surface electrons into the bulk of the system, leading to the broadening of the electronic resonances. The broadening at 4.6 K is generally smaller in Ag compared to Cu [81]. However, the adatoms placed on Cu, especially CO molecules, are more stable and easier to move on Cu than Ag, making Cu the metallic reference system for creating artificial lattice structures [93]. This combination has been employed for designing lattices with translational invariance [102, 243, 242, 129, 92, 159].
The molecular self-assembly approach deposits molecular precursors on a metallic surface; after annealing, the precursors rearrange themselves in structures that can be more complex than the original precursors, either by following a specific pattern intrinsic to the precursor or because of some property of the substrate. Successively, the structure is analyzed either by STM or by an atomic force microscope [205].
While the before-mentioned approaches are, at present, limited to single-particle physics, twisted two-dimensional materials provide tunable lattices that showcase strongly correlated physics. A prominent example is twisted bilayer graphene exhibiting unconventional superconductivity at a magic twist angle [43]. Quite generically, twisting 2D lattices leads to a Moiré superlattice determined by the twist angle, yielding the basis for tunable electronic materials that can be used for quantum simulation [131]. One advantage of this platform is that adding more layers or combining different materials (e.g. layers of transition metal dichalcogenides) permits the exploration of quantum simulators with diverse properties.
@@ II.3.1 Quasicrystals and fractals
Twisted 2D layers.
Stacking and twisting layers of 2D materials is a natural way to obtain large-scale quasicrystalline structures. A quasiperiodic lattice with 12-fold rotational symmetry is obtained from twisting two hexagonal lattice structures by an angle of 30∘ relative to each other [249]. This quasicrystal has been realized with twisted graphene bilayers [5, 306], and twisted transition metal dichalcogenides [161]. It has then been proposed that adding a further twisted layer would permit the combination of the flat-band physics at small twist angles, and the quasiperiodic structure appearing at larger angles can be combined [271]. Such quasicrystal flat bands have then been observed in twisted trilayer graphene [107], providing a feasible setup for strongly correlated phases of quasicrystals.
Circuit QED.
Superconducting qubits can implement Bose-Hubbard models with programmable on-site chemical potential and programmable hopping terms. This has been used in Roushan et al. [219], Li et al. [156] to build short chains (up to 10 sites) with quasiperiodic potential and/or tunneling, implementing 1D quasicrystals of bosons.
@@ II.3.3 Higher dimensional physics
Classical electric circuits are also a versatile platform for implementing higher-dimensional physics. The two-dimensional array of capacitors and inductors can be easily extended to realize periodic 3D and 4D circuits on a breadboard [315]. The idea is to decompose an D structure into a set of connected D circuit boards; see Fig. 8.
@@
II.4 Summary
As demonstrated by the above sections, for each of the exotic geometries considered in this colloquium, various possible implementations exist that use atoms, photons, or electrons. Despite the versatility of each platform, they have strengths and weaknesses, which we outline in Tab. 2. Thereby, one should bear in mind that, at least to some extent, these categories are rather subjective, and a nuisance in one context can be an opportunity in another one.